OpenStudy (anonymous):

WILL MEDAL PLEASE ANSWER!!! Which data set has the widest spread? data set A: 1, 1, 2, 1, 2, 1 data set B: 2, 2, 4, 4, 3, 3 data set C: 1, 1, 2, 2, 1, 5 data set D: 3, 5, 5, 5, 5, 4

OpenStudy (anonymous):

Please answer

OpenStudy (kira_yamato):

From the way the question is phrased, I think you are referring to standard deviation (the sample has N items) (I'll denote with σ). You have to use the formula \[\sigma = \frac{1}{N}\sum_{all \space x} (x_i - \bar{x})\] Your answer should be the sample that gives you the largest value for σ

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Witch one is it @Kira_Yamato

OpenStudy (anonymous):

It is still confusing

OpenStudy (anonymous):

BTW it is Statisics

OpenStudy (anonymous):

Rickey's aunt has mailed him a few packets of collectible stamps. Rickey knows that each packet contains 5 unique stamps, because he already has 4 such packets. If his aunt has sent him p packets of stamps, which expression gives the total number of stamps Rickey will have after receiving his gift? 5p + 4 4p + 5 5(p + 4) p(4 + 5)

OpenStudy (kira_yamato):

You have to first find the mean for each sample. Then take each number minus mean and square that answer. Sum everything up you have what is known as variance. SD is just taking sqrt of the variance

OpenStudy (anonymous):

and equals up to...

OpenStudy (kira_yamato):

Sorry after taking sqrt you need to divide by number of items in sample before you get SD So SD(A) = sqrt[4 x (1-8/6)^2 + 2 x (2-8/6)^2]/6 = sqrt(8/9)/6 = √2 / 9 ≈ 0.15713 I'll leave the rest to you to calculate. It's similar